a rectangle with a perimeter of 40 centimeters is 5 centimeters longer than twice its width. find the length of each side, then find the area

Answer by checkley77(12844)

L=2W+5
2L+2W=40
2(2W+5)+2W=40
4W+10+2W=40
6W=40-10
6W=30
W=30/6
W=5 ANS. FOR THE WIDTH.
L=2*5+5
L=10+5=15 ANS. FOR THE LENGTH.
PROOF:
2*15+2*5=40
30+10=40
40=40